### v1

@@
These authors contributed equally to this work.
[…]
Characterizing the Reasoning Abilities of Large Language Models via 3-SAT Phase Transitions// The Ability to Reason: A Characterization of LLMs via 3-SAT
[…]
keywords:
keyword1, Keyword2, Keyword3, Keyword4
1 Introduction
The success and versatility of Large Language Models (LLMs) has sparked a widespread interest in their reasoning abilities. (Luc) Let’s mix that sentence There are different views on what reasoning in this context means, whether it is oriented towards common-sense reasoning (REFS) or towards logical or deductive reasoning. We will adhere to Leon Bottou’s definition, which states that reasoning is “[the] algebraic manipulation of previously acquired knowledge in order to answer a new question”. Recent studies suggest that LLMs are inherently capable of zero-shot reasoning [llms_zero_shot_reasoners], that is, performing multi-step inference processes in previously unseen situations. Combined with smart prompting mechanisms that encourage LLMs to think step-by-step, the reasoning abilities of LLMs can be enhanced [chain_of_thought, leasttomost, react, adapt]. Demonstrations of the reasoning abilities of LLMs include but are not limited to planning [llms_zero_shot_planners, saycan, progprompt, codeaspolicies, innermonologue, saycanpay], theorem proving [jiang2023draft, natural_prover, verifier_guided_proof], search and optimization [llms_as_optimizers, llmatic, fun_search], self-reflection [tree_of_thoughts, selfrefine, reflexion], API calling [hugginggpt, toolformer], etc.
On the other hand, there is a significant body of research that presents a critical view of the reasoning abilities of LLMs. These studies highlight cases where LLMs struggle to provide the correct answers. For instance, LLMs may exhibit limitations in consistent logical reasoning [gpt4_cant_reason, llms_are_greedy_reasoners], effective planning [llms_cant_plan], and conducting accurate self-assessment of their outputs [gpt_graph_coloring]. These contrasting perspectives create an ambiguous picture about the true reasoning abilities of LLMs.
According to Leon Bottou [13] reasoning is “[the] algebraic manipulation of previously acquired knowledge in order to answer a new question”.
2 3-SAT problems
To clarify this picture, we evaluate the performance of LLMs on solving 3-SAT problems. 3-SAT is one of the most fundamental problems in computer science and reasoning. The 3-SAT problem is introduced via the SAT-game in BOX 1 (and state them also in implicative form).
It is the prototypical NP-complete problem lying at the basis of computational complexity. To date it is unknown whether there exist efficient (that is, polynomial time) algorithms for NP-complete problems. Numerous important computational problems are NP-complete and can be efficiently reduced to 3-SAT. Furthermore, any polynomial time algorithm for 3-SAT would result in an efficient algorithm for these other NP-complete problems. One other such NP-complete problem is $k$-SAT, which generalizes 3-SAT by lifting the restriction of 3 literals per clause. SAT-solvers are frequently used to solve reasoning problems. First, $k$-SAT problems are constraint satisfaction problems (over boolean variables). Second, SAT-solvers are used to prove that a sentence $s$ is logically entailed by a set of clauses $S$ (notation $S\models s$) using the principle of reductio ad absurdum, which corresponds to adding the negation of the rule $NOT~{}s$ to the set of clauses $S$ and showing that $S~{}AND~{}(NOT~{}s)$ is unsatisfiable. Third, various other reasoning problems in planning and constraint satisfaction can be solved by reducing them to SAT-solvers.
$k$-SAT is well-studied, both theoretically and empirically. The later is especially relevant to obtain insight into the ability of LLMs to reason. It is well-known that there are easy and hard problem instances in 3-SAT (SELMAN et al). Considering the 3-SAT game with a fixed number of variables in Box 1, the number of solutions decreases with the number of clauses. Figure 1 depicts the probability that there exists a satisfying assignment for a particular problem as a function of $N/V$ where $V$ is the number of variables and $N$ the number of rules. It also shows the the phase transition that divides the space into the typically easy - hard - easy regions. When there are only few rules, it is very likely that there exists a solution and it is easy to find it. When there are very many rules, it is unlikely that a solution exists, and close to the phase transition one finds the really hard problems. By analogy to the use of these insights to assess the difficulty of SAT and the performance of solvers, we will use them to empirically assess the reasoning abilities of LLMs.
(Luc) There are also the theoretical arguments (about complexity) that we should write a sentence about. (Gabriele) I’m not entirely sure what theoretical arguments you are referring to..? If you clarify I can try to write something about it. Also, I think something is missing from the story to convince the readers that 3-SAT is a good candidate to do this reasoning-capability study for LLMs. Maybe I would stress a bit more the importance of the NP class. Possibly, you were referring to this?
3 Evaluating the reasoning ability of LLMs on 3-SAT
To evaluate the extent to which LLMs can solve 3-SAT problems, we consider both the decision and the search variant of the 3-SAT problem. In the 3-SAT decision problem, the solver is given a 3-SAT formula and has to return YES when there exists a satisfying assignment and NO when there is no such assignment. In the search version of 3-SAT, the solver must return a satisfying assignment if such an assignment exists, and NO otherwise.
We are seeking answers to the following experimental questions:
(1) Which type of 3-SAT problems can LLMs solve?
(2) Can LLMs use a solver to enhance their reasoning ability?
(3) Are there differences in the solving ability of state-of-the-art LLMs?
We use GPT-4 as the reference LLM, given its superior performance in existing benchmarks [^1] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard. Subsequently, to answer question (3) we extend our analysis to include comparisons with various other state-of-the-art LLMs.
Furthermore, as the type of input and the prompting strongly influence the results obtained with LLMs, we consider two types of input: (i) SAT-Menu uses a menu-selection scenario to which the CNF formula has been mapped (see Figure X), and (ii) SAT-CNF uses the CNF formula directly as input (see Appendix Table C), and we explore various prompting techniques incorporating in-context learning [llms_few_shot_learners] and chain-of-thought (CoT) [chain_of_thought] prompts. Moreover, we always instruct the LLMs to think step-by-step [llms_zero_shot_reasoners].
Which type of 3-SAT problems can LLMs solve?
  • 3-SAT Decision Problem: As shown in the confusion matrix (Figure 3 [Right]), GPT-4 shows proficiency in identifying unSAT scenarios, though it occasionally misclassifies SAT problems as unSAT, particularly within the Hard region of $\alpha$.
  • 3-SAT Search Problem: Investigating GPT-4’s handling of the 3-SAT search problem reveals a phase transition pattern similar to that of conventional solvers, characterized by Easy-Hard-Easy phases (Figure 2). While GPT-4 demonstrates competent performance in the Easy region, its accuracy significantly diminishes to $\approx 10\%$ in the Hard region. This pattern underscores the model’s variable effectiveness across the phase transition – how GPT-4 is significantly better at reasoning through inherently easy problems than with hard problems.
  • Size of solution space: In Figure 3[Left], we also plot the performance against the satisfiability ratio, given as model $\frac{\text{model count}}{2^{n}}$. Here, model count gives the number of feasible solutions and $n$ is the number of variables. Thus the satisfiability ratio denotes the probability that a randomly selected variable assignment satisfies the given 3-SAT theory. The higher this probability the more likely it is that GPT-4 finds a solution.
  • Impact of Input Type on Reasoning: We observed no change in performance with direct SAT-CNF inputs over SAT-Menu (Figure 9). Thus, the phase transitions remain the same regardless of the input.
  • Impact of Prompt Engineering: Prior works have shown that in-context learning can enhance LLM performance [llms_few_shot_learners]. We explored this with 3-shot learning using three input/output (I/O) examples in the prompt. These examples were randomly selected from a set of accurate solutions generated by the LLM and manually checked for consistency between solutions and their explanations. Each I/O example included both the input preferences and the output solution along with its explanation. From Figure 4 [Left], we saw performance gains in the initial Easy-Hard phase but observed a decrease in the subsequent Easy phase. (Rishi) Explain …
(Luc) What is our explanation / conclusion for these experiments ? Our findings reveal that irrespective of prompt engineering techniques, GPT-4 struggles in reasoning, especially in the Hard region. Even within the Easy region, GPT-4’s performance drops if the satisfiability ratio is low.
(Luc) The differences between the decision problem and solving the SAT search problem is not always clear in the graphs.
Can LLMs use a solver to enhance their reasoning ability?
Here we feed the LLM with the SAT-Menu input and prompt it to map it to a CNF formula. The resulting approach, termed SAT-Translate, thus uses an LLM to translate a natural language task into a solver-compliant SAT formula and then applies a SAT solver for deriving a solution (see Appendix Box D). This underscores the synergy between the LLMs’ capabilities and the solvers’ computational precision [word_to_world_models, dinu2024symbolicai, llmp, linc, xie2023translating, thor_theorem_proving, mathematical_conjecturing_with_llms]. As shown in Figure 4 [Right], the performance is almost perfect. This success is attributed to the relatively lower complexity of translation tasks compared to the NP-complete nature of solving SAT-problems.
Are there differences in the solving ability of state-of-the-art LLMs?
We have conducted comparative experiments with some of the best and most advanced LLMs. This includes open-source (Llama 2 70B (chat-hf) [meta_llama2], Mixtral [mistral_mixtral]) as well as proprietary models (GPT-4 (Turbo) [openai_gpt4], GPT-3.5 (Turbo) [openai_gpt3_5], Gemini Pro [google_gemini], PaLM 2 (text-bison) [google_palm]). In these evaluations, GPT-4 emerges as a notable exception, outshining other LLMs across various metrics (cf. the accompanying material).
We identify both common and divergent performance patterns.
Common Performance Trends:
  • Size of solution space: Across LLMs, performance correlates with solution space size (Figure 8).
  • SAT-Menu vs. SAT-CNF: In general, LLM performance does not change with input type (Figure 9).
  • SAT-Translate: Integration of LLMs with solvers is more effective (Figure 9).
Divergent Performance Trends:
  • Phase Transitions: As shown in Appendix Figure 7 [Left], GPT-4 uniquely reflects solver-like phase transitions. In contrast, other LLMs exhibit a marked decline in performance in the high $\alpha$ region, displaying an Easy-Hard-Hard pattern. This divergence primarily stems from their failure to identify unSAT problems, which predominate at higher $\alpha$ values (see Appendix Figure 7 [Right]).
  • In-context improvements: While GPT-4 shows improvements, other LLMs show no performance change with in-context learning (Figure 10).
3.1 Can LLMs use a solver to enhance their reasoning ability?
BOX 2: show the inputs for a LLM.
Figure 1: The SAT Game is a gamification of the SAT problem. In the SAT game, one must find a valuation of Boolean variables (true/false) such that a Boolean CNF formula evaluates to true. A line in the game actually represents a clause, which is a disjunction of literals (variables or their negation connected by logical OR). A clause evaluates to true as soon as one of its literals is true. Since clauses are conjunctively connected (by a logical AND), all clauses must evaluate to true so that the formula be true.
4 Related Works
We group the papers into three categories – Papers that perform an in-depth exploration of reasoning abilities; papers that perform an empirical evaluation of reasoning abilities on benchmark datasets and tasks; and finally papers that propose methods to improve reasoning.
Intensive Studies
Recent works have examined how transformer architecture, training, and inference paradigms influence the reasoning abilities of LMs. Training analysis shows how LMs like Bert can fit on statistical features, instead of emulating correct reasoning, to achieve high performance [paradox_of_learning_to_reason]. This is also linked to the Clever Hans Cheat in teacher-forcing training where LLMs identify shortcuts that bypass true reasoning [pitfalls_of_next_token_prediction]. In terms of inference, the autoregressive nature of LLMs make them prone to snowballing errors over time [dziri2023faith]. A more systematic approach is to directly evaluate the predicted chains of thought, wherein it has been revealed how LLMs are prone to generating unfaithful and biased explanations [llms_dont_say_what_they_think]. Moreover, LLMs reason greedily, failing in contexts with multiple valid reasoning steps [llms_are_greedy_reasoners].
On the architectural side, recent research has shown that the transformer layer is incapable of function composition for larger domains [limits_of_transformer_architectures, dziri2023faith]. Moreover, it has been shown that the computation of transformers lies in the complexity class log-uniform TC0, which is more restrictive than the logarithmic memory class L [log_precision_transformers]. Given that the $2-$SAT problem is NL-complete (i.e., the class of non-deterministic logarithmic space algorithms), multi-layer transformers cannot solve $2-$SAT instances, unless L$=$NL [limits_of_transformer_architectures].
In this work, we delve into reasoning [^2] We specifically focus on search-based reasoning while excluding domains like commonsense reasoning, an emergent ability of LLMs, shown to manifest with scale [wei2022emergent, llms_few_shot_learners, bigbench]. However, existing works have also brought this claim to question, arguing if the emergent abilities are but a mirage stemming from choice of metric rather than genuine model improvements [llms_emergent_abilities_a_mirage]. Our research aligns closely with NPHardEval [nphardeval], which examines reasoning across various computational complexity classes, but we additionally explore phase transition characteristics and investigate how the performance varies with inherent hardness of problems. Moreover, our evaluation centers on the 3-SAT, a classic NP-complete problem known for its distinct phase transition characteristics, providing a clear framework for analyzing reasoning depth.
Empirical Studies
Recent studies suggest that Large Language Models (LLMs) are inherently capable of zero-shot reasoning [llms_zero_shot_reasoners]. Other reasoning abilities of LLMs include but are not limited to, planning [llms_zero_shot_planners, progprompt, codeaspolicies, innermonologue], theorem proving [jiang2023draft, natural_prover, verifier_guided_proof], search and optimization [llms_as_optimizers, llmatic, fun_search], self-reflection [tree_of_thoughts, selfrefine, reflexion], API calling [hugginggpt, toolformer].
Contrastingly, there is a significant body of research that presents a more critical view of LLMs’ capabilities. These studies highlight the struggles of LLMs in areas where they are presumed to excel. For instance, they may exhibit limitations in logical reasoning [gpt4_cant_reason, llms_are_greedy_reasoners], planning [llms_cant_plan], and conducting accurate self-assessment of their outputs [gpt_graph_coloring]. The contrasting perspectives from various studies create an ambiguous picture of LLMs’ true abilities. This disparity prompts a fundamental question: Are LLMs truly capable of reasoning, and if so, to what extent? In this work, we study the phase transition characteristics of LLMs on 3-SAT problems. This enables us to ground the reasoning abilities of LLMs into a more formal framework and helps us study the performance changes across varying problem hardness, size of solution space, and type of inputs.
Strategies for Improving Reasoning
Previous research has demonstrated that the reasoning capabilities of LLMs can be improved through the use of intelligent prompting techniques that guide the models to process information step-by-step [chain_of_thought, leasttomost, react, adapt]. Additionally, the integration of LLMs with Neuro-Symbolic frameworks [deraedt2020statistical] leverages the systematic approach of symbolic solvers alongside the strengths of LLMs. This includes reasoning [word_to_world_models, dinu2024symbolicai], planning [llmp, linc, xie2023translating, llms_as_world_models_plus_planners], automated conjecturing [mathematical_conjecturing_with_llms], automated theorem proving [thor_theorem_proving], and so on.
Another innovative direction involves augmenting LLMs with external tools [toolformer, hugginggpt] and verifiers [verify_step_by_step, saycan, saycanpay, llm_modulo_frameworks] to improve its performance.
Figure 2: Surface plot illustrates the phase transition comparison between GPT-4 (left) and solver MiniSAT (right). It can be observed that GPT-4 mirrors the (inverted) Easy-Hard-Easy characteristics of the solver wherein the Hard region is defined by $\alpha\in[3,5]$. Note that the LLM phase transitions are on accuracy while the solver phase transitions are on time taken to solve. The LLM setup is SAT-CNF with 0-shot prompting. (Luc) Why do not we just plot the probability that a solution exists for a random problem - time taken does not matter, that is an indirect effect.
Figure 3: [Left] Figure demonstrates how an increase in satisfiability ratio leads to improvements in accuracy. Specifically, it highlights that GPT-4 is more likely to produce a satisfying solution as the probability of identifying such a solution increases for the specified problem. (Luc) Observe that we only include satisiable problems whose satisfiability ratio is strictly larger than 0. [Right] Confusion matrix shows the accuracy of GPT-4 on the 3-SAT decision problem. The cell annotations reflecting classification accuracy, normalized over the true count (columns) to account for the imbalance between SAT and unSAT instances. The setup is SAT-Menu with 0-shot prompting.
Figure 4: [Left] Comparison of 0-shot and 3-shot performance on SAT-Menu showing GPT-4 overall performance improves with in-context learning. [Right] GPT-4 performance on the SAT-Translate problem, highlighting the benefits of augmenting LLMs with solvers.
5 Results
(1) Can GPT-4 effectively reason through diverse problem complexities, or is it limited to inherently simpler tasks? To answer this question, we perform the following experiments:
  • 3-SAT Optimization Problem: Investigating GPT-4’s handling of the 3-SAT optimization problem reveals a phase transition pattern similar to that of conventional solvers, characterized by Easy-Hard-Easy phases (Figure 2). While GPT-4 demonstrates competent performance in the Easy region, its accuracy significantly diminishes to $\approx 10\%$ in the Hard region. This pattern underscores the model’s variable effectiveness across different problem complexities – how LLMs are significantly better at reasoning through inherently easy problems than they are with hard problems. Note, that for LLMs, we plot the 3-SAT optimization accuracy, instead of the solution time used for solvers.
  • Size of solution space: We also plot the performance against the satisfiability ratio which is given as model $\frac{\text{model count}}{2^{n}}$ in Figure 3 [Left]. Here, model count gives the number of feasible solutions and $n$ is the number of variables. This signifies the number of feasible solutions in the entire range of solutions with $n$ variables. We observe that a larger solution space correlates with improved performance. I (Luc) can we conned sat. ratio to phase transition ?
    (Luc) I do not understand this formulation
  • 3-SAT Decision Problem: As shown in the confusion matrix (Figure 3 [Right]), GPT-4 shows proficiency in identifying unSAT scenarios, though it occasionally misclassifies SAT problems as unSAT, particularly within the Hard region of $\alpha$.
  • Impact of Input Type on Reasoning: We observed no change in performance with direct SAT-CNF inputs over SAT-Menu (Figure 9). Thus, the phase transitions remain the same regardless of the input.
  • Impact of Prompt Engineering: Prior works have shown that in-context learning can enhance LLM performance [llms_few_shot_learners]. We explored this with 3-shot learning using three I/O examples in the prompt. These examples were randomly picked from a set of accurate solutions generated by the LLM and manually checked for consistency between solutions and their explanations. Each I/O example included both the input preferences and the output solution along with its explanation. From Figure 4 [Left], we saw performance gains in the initial Easy-Hard phase but observed a decrease in the subsequent Easy phase. (Rishi) Explain …
(2) Does integrating a solver with GPT-4 enhance its problem-solving capabilities? This approach, termed SAT-Translate, involves using an LLM to translate a natural language task into a solver-compliant formula and then applying a solver for solution derivation (see Appendix Box D). This underscores the synergy between LLMs’ capabilities and solvers’ computational precision [word_to_world_models, dinu2024symbolicai, llmp, linc, xie2023translating, thor_theorem_proving, mathematical_conjecturing_with_llms]. Here we feed the LLM with the SAT-Menu input and prompt it to map it to a CNF formula. As shown in Figure 4 [Right], the performance is almost perfect. This success is attributed to the relatively lower complexity of translation tasks compared to the NP-complete nature of solving tasks.
(3) Is there any variability in the general reasoning abilities of state-of-the-art LLMs? In evaluating state-of-the-art LLMs, GPT-4 emerges as a notable exception, outshining other LLMs across various metrics. We identify both common and divergent performance patterns.
Common Performance Trends:
  • Size of solution space: Across LLMs, performance correlates with solution space size (Figure 8).
  • SAT-Menu vs. SAT-CNF: In general, LLM performance does not change with input type (Figure 9).
  • SAT-Translate: Integration of LLMs with solvers is more effective (Figure 9).
Divergent Performance Trends:
  • Phase Transitions: As shown in Appendix Figure 7 [Left], GPT-4 uniquely reflects solver-like phase transitions. In contrast, other LLMs exhibit a marked decline in performance in the high $\alpha$ region, displaying an Easy-Hard-Hard pattern. This divergence primarily stems from their failure to identify unSAT problems, which predominate at higher $\alpha$ values (see Appendix Figure 7 [Right]).
  • In-context improvements: While GPT-4 shows improvements, other LLMs show no performance change with in-context learning (Figure 10).
6 Introduction
Large Language Models (LLMs) have rapidly captured the attention of the AI research community with their exceptional natural language completion capabilities. Trained on web-scale language corpora, these models have demonstrated the ability to generate seemingly valuable completions across a wide range of topics. Their versatility has sparked widespread interest and debate regarding their potential role in advancing fields such as mathematics, coding, vision, medicine, law, psychology, and beyond – and how some of these models could reasonably be viewed as an early (yet still incomplete) version of an artificial general intelligence (AGI) system [sparks_of_agi].
Yet, from the perspective of computational theory, we are reminded that numerous practical computational tasks are surprisingly difficult to solve even with the most advanced computing solutions. Such problems, found for example in planning, scheduling, machine learning, hardware design, and computational biology, generally belong to the class of NP-complete problems. NP stands for ‘nondeterministic polynomial time’, a class of computational tasks consisting of problems for which a potential solution can be checked efficiently for correctness, yet finding such a solution appears to require exponential time in the worst case. The completeness property of NP-complete problems means that if an efficient algorithm for solving just one of these problems could be found, one would immediately have an efficient algorithm for all NP-complete problems.
Boolean satisfiability problem (SAT), and more specifically, k-SAT is an an archetypal NP-complete problem where given a Boolean formula composed of variables and logical operators AND, OR, and NOT, one determines if there is a way to assign truth values (TRUE or FALSE) to the variables such that the entire formula evaluates to true. It involves searching through the space of all possible assignments of truth values to variables to find an assignment that “satisfies” the formula. This parallels many reasoning tasks in AI like logical reasoning and constraint satisfaction.
The concept of phase transition in the SAT problem is crucial. It refers to a sharp change in the probability that a randomly generated $k$-SAT instance is satisfiable as a parameter of the problem (usually the ratio of clauses to variables) crosses a critical threshold. The hardest instances of k-SAT problems are typically found around this threshold. Algorithms that perform well on these instances are likely to be robust and efficient in reasoning in a variety of settings. This not only helps in gaining insights into the performance limits and efficiency of AI algorithms but also aids in the design of more effective AI systems – if an AI system can handle problems in this region effectively, it is a strong indicator of its overall capability.
@@ 6 Introduction
Figure 5: Dataset Statistics. [Left] The inner segment of the nested pie chart illustrates how variables are distributed, whereas the outer segment displays the breakdown of SAT and unSAT instances for each variable. arises from the varying $\alpha$ values assigned to each variable, detailed in Table 1. [Right] Figures depict clauses and alpha distribution across SAT and unSAT instances, highlighting that unSAT problems typically feature more clauses and higher alpha values.
Appendix A Dataset
A.1 Generation
In this section, we outline the dataset employed for our study and the methodology behind its generation.
  • We generated two two distinct datasets: the first with $k=2$ targeting polynomial problems (2-SAT), and the second with $k=3$ for NP-complete problems (3-SAT).
  • We employed $\alpha=m/n$ as a parameter to guide dataset generation across the $\alpha$ range of $[1,11]$. Here, $m$ is the number of clauses, and $n$ is the number of variables.
  • For $\alpha$ values within the $(6,11]$ interval, we incremented $\alpha$ by $1$. For the $[1,6]$ interval, which contains the most “interesting problems,” we aimed for finer granularity by choosing the smallest possible $\alpha$ increment. This increment ensures that, given the number of variables $n$, we obtain an integer number of clauses $m$. For example, with $n=3$, the minimum increment is $\alpha_{inc}=1$, and for $n=4$, it is $\alpha_{inc}=0.25$.
  • We generated formulas with $n\in[1,10]$ variables. For each $\alpha_{inc}$ we generated $300$ formulas.
  • In the dataset with $k=2$, we applied the same methodology but generated formulas within an $\alpha$ range of $[1,10]$, always using the smallest feasible increment for each $\alpha$. All other parameters remained consistent.
  • Each instance is labeled either as satisfiable or unsatisfiable using MiniSAT v2.2 [een2003extensible].
  • Each instance in the dataset is annotated with model count (i.e. number of feasible solutions for a formula). We used D4 [lagniez2017improved] for model counting.
Table 1: Table shows range of alpha value for each $n$ (i.e. number of variables) in the generated dataset. We generate $300$ formulas per $\alpha$ value.

### v2

@@ 6 Discussion
Even though the basic random 3-SAT phase transition provides a rigorous framework for studying the reasoning capabilities of LLMs, we need to point out that there exists many more results about 3-SAT that might be considered as well. For instance, it has been found that certain aspects of the hardness of 3-SAT instances cannot be explained using the easy-hard-easy regimes induced by the phase transition (Mu & Hoos, 2015). Nevertheless, the phase transition seems to be adequate in our settings as demonstrated by the drop in performance of LLMs in the vicinity of $\alpha_{c}$. It should, however, be noted that 3-SAT instances encountered in practice often fall into the easy regions. Therefore, decision problems of interest do not usually exhibit worst-case complexity.
7 Conclusion
A superficial analysis of the reasoning capabilities of LLMs suggests they possess strong and complex reasoning capabilities – a common fallacy (Bubeck et al., 2023). However, our detailed experimental analysis indicates that what appears as reasoning capabilities could be a mirage, with LLMs predominantly exploiting statistical features present in the data and would explain why LLMs have been termed “statistical parrots” (Bender et al., 2021).
This is not to say that LLMs lack value – quite the opposite. LLMs are highly effective at translating problems into a formal language and then passing these problems on to solvers. This utility is demonstrated by the relatively superior performance of SAT-Translate over SAT-Menu and SAT-CNF. In a more general context, it would require that the class of the problem is recognized correctly – here 3-SAT – and that the right solvers are available.
Furthermore, LLMs can serve as valuable knowledge bases, leveraging their extensive common sense and world knowledge through natural language queries to guide search processes. This can be utilized by either solvers guiding LLMs (Hazra et al., 2024b) or vice versa, where LLMs assist solvers (Wong et al., 2023).
Reproducibility
We provide comprehensive instructions to ensure the reproducibility of the results. The dataset generation process and statistics are outlined in § 4.2 and Appendix A, the LLM configurations are listed in Appendix Table 2, and detailed prompts shown in Appendix C. Moreover, we also discuss limiting/failure cases in Appendix B and how they can be mitigated. We intend to open-source our code shortly.
Acknowledgments
This work was supported by the Wallenberg AI Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, by the EU H2020 ICT48 project “TAILOR” under contract #952215, and the KU Leuven Research Fund (C14/18/062). The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government. We also thank Holger Hoos, Heikki Mannila, Paolo Frasconi, Pascal Van Hentenryck, Ross King, Giuseppe Marra, Pieter Delobelle, Hendrik Blockeel, and Raffaele Marino for their valuable feedback.
@@ Acknowledgments
The appendix is organized as follows. Dataset Statistics § A, Failure Cases of LLMs § B, and Full Prompts § C.
Figure 6: Dataset Statistics. Figures depict clauses $m$ (left) and $\alpha$ (right) distribution across SAT and unSAT instances, highlighting that unSAT problems typically feature more clauses and higher $\alpha$ values.
Appendix A Dataset Statistics
Table 1 lists all values of the range of $\alpha$ corresponding to each $n$. For $\alpha$ values within the $(6,11]$ interval, we incremented $\alpha$ by $1$. For the $[1,6]$ interval, which contains the most “interesting problems,” we aimed for finer granularity by choosing the smallest possible $\alpha$ increment. This increment ensures that, given the number of variables $n$, we obtain an integer number of clauses $m$. For example, with $n=3$, the minimum increment is $\alpha_{inc}=1$, and for $n=4$, it is $\alpha_{inc}=0.25$. For each $\alpha_{inc}$ we generated $300$ formulas.
The distribution of formulas according to the number of variables is detailed as follows: 3,000 formulas with 3 variables, 7,500 with 4 variables, 9,000 with 5 variables, 4,500 with 6 variables, 3,000 with 7 variables, 13,500 with 8 variables, 3,000 with 9 variables, and 16,500 with 10 variables.
As shown in Figure 6, the range of clauses ($m$) varies in [5, 90], and the alpha ($\alpha=m/n$) ranges in [1.1, 11.0]. For unSAT instances, the clauses range from 9 to 110, with the alpha varying in [2.60, 11.0]. Across the entire dataset, the average number of variables ($n$) is 7.2, the average number of clauses ($m$) is 33, and the mean $\alpha$ is 4.7. This diverse dataset provides a broad spectrum for analyzing the impact of variable and clause distribution on formula satisfiability.
Table 1: Table shows the range of alpha value for each $n$ (i.e. number of variables) in the generated dataset. We generate $300$ formulas per $\alpha$ value.